OpenStudy (anonymous):

i really need help.... if one solution of the equation x^2+ m = 0 is 5i, find the value of m.

OpenStudy (anonymous):

m = 5

OpenStudy (mertsj):

replace x with 5i and solve for m

OpenStudy (anonymous):

X^2=-5

OpenStudy (anonymous):

square root both sides

OpenStudy (mertsj):

@salparadise64 Might want to rethink your advice.

OpenStudy (anonymous):

x= 5i

OpenStudy (anonymous):

@salparadise64 @Mertsj wouldnt it be -25?

OpenStudy (anonymous):

bro, i know what im doing

OpenStudy (mertsj):

\[(5i)^2+m=0\] \[-25+m=0\] \[m=25\]

OpenStudy (anonymous):

i have a bachelors in applied mathematics

OpenStudy (mertsj):

Congratulations.

OpenStudy (anonymous):

so shove off poser

OpenStudy (mertsj):

\[x^2+25=0\] has one solution of 5i and another solution of -5i

OpenStudy (mertsj):

Yes, @salparadise64 , x = 5i. That was given. Your job was to solve for m

OpenStudy (anonymous):

extraneous roots bro bro. you can't go around introducing numerical values like that. where did you get your bachelors in applied mathematics from?

OpenStudy (anonymous):

oh, is the answer "nowhere" or "i don't have one" ? then i suggest you listen to a real graduate.

OpenStudy (anonymous):

@salparadise64 @Mertsj lmao its just math problem guys but thanks alot of your help

OpenStudy (mertsj):

@salparadise64 The problem said to find the value of m in the equation x^2+m=0 if x = 5i What did you get for m?

OpenStudy (anonymous):

no problem bro, anytime you need help just message directly to avoid jokers like @Mertsj

OpenStudy (mertsj):

@salparadise64 What did you get for m?